# Tight frames over the quaternions and equiangular lines

### v2

@@ 1 Introduction
There exists 2d2d equiangular lines in d, for each $\hbox{\em There exists $2d^{2}-d$ equiangular lines in $\mathbb{H}^{d}$, for each $d$},$
should hold. However, calculations of [CKM16], suggests that this fails for $d=4$, and the analogous situation for the octonians is much worse. Thus, it seems that equiangular lines in $\mathbb{C}^{d}$ may be a high point for satisfying the estimates on the maximal number of equiangular lines, with real and quaternionic equiangular lines rarely meeting the bound (“filling up all the space”) due to algebraic limitations of the field involved, i.e., $\mathbb{R}$ not being algebraically closed and $\mathbb{H}$ not being commutative. Still, there is much interest in the maximal sets sets of equiangular lines in $\mathbb{H}^{d}$, and for those in $\mathbb{R}^{d}$ (which have been studied for over half a century). Therefore, I present the following conjecture, which can play the role of Zauner’s conjecture for the theory of quaternionic equiangular lines:
@@ 1.1 Inner products over the quaternions
Since the multiplication is not commutative, we must distinguish between left and right vector spaces (modules) over $\mathbb{H}$. Since we wish to appropriate much of matrix theory, we take our vector spaces to be right $\mathbb{H}$-vector spaces. Thus $\mathbb{H}$-linear maps have the form
@@
where order of multiplication in $a_{j\ell}b_{\ell k}$ cannot be reversed. For those who may have noticed, I apologise for using $j$ and $k$ above as indices for matrix entries, and elsewhere as quaternian units (as is often done with the complex unit $i$).
@@
$\overline{ab}=\overline{b}\,\overline{a},\qquad a,b\in\mathbb{H}.$
@@
$\langle v,w\rangle:=\sum_{j}\overline{v_{j}}w_{j},\qquad v,w\in\mathbb{H}^{d}.$ (1.2)
@@
In light of the noncommutativity of the quaternians, we note that scalars come outside an inner product (as we have defined it) as follows
@@
$\langle v,w\rangle=0\hskip 10.00002pt\Longleftrightarrow\hskip 10.00002pt\langle w,v\rangle=\overline{\langle v,w\rangle}=0.$
@@
If $T$ and $T^{*}$ are represented as matrices $[T]$ and $[T^{*}]$ with respect to orthonormal bases $(v_{j})$ and $(w_{k})$, so that $v=\sum_{j}v_{j}\langle v,v_{j}\rangle$, $\forall v\in{\cal V}$, and $w=\sum_{k}w_{k}\langle w,w_{k}\rangle$, $\forall w\in{\cal W}$, then
$[T]_{jk}=\langle Tv_{k},w_{j}\rangle=\langle v_{k},T^{*}w_{j}\rangle=\overline{\langle T^{*}w_{j},v_{k}\rangle}=\overline{[T^{*}]_{kj}},$
and hence the matrix $[T^{*}]$ is the conjugate transpose of the matrix $[T]$. For this reason, it is often assumed that the inner product is the standard inner product on $\mathbb{H}^{d}$, and all calculations are done with matrices, with $A^{*}$ defined to be the conjugate transpose (or Hermitian transpose) of the matrix $A$, as is the case in [Rod14]. The adjoint (or Hermitian transpose) satisfies some (but not all) of the usual properties, including
@@
$(A^{*})^{-1}=(A^{-1})^{*}\quad\hbox{(for $A$ invertible)}.$) .
It can be shown that if $AB=I$ (for matrices over $\mathbb{H}$) then $BA=I$, and so a right inverse exists for $A$ if and only if a left inverse exists, and these inverses are equal (and denoted by $A^{-1}$).
@@
[…]
Left multiplication of $\mathbb{H}^{d}$ by $\beta$ defines an $\mathbb{H}$-linear map $L_{\beta}:\mathbb{H}^{d}\to\mathbb{H}^{d}:v\mapsto\beta v$, but this is dependent on a choice of basis: it is the linear map which maps $e_{j}\mapsto e_{j}\beta$, i.e., the linear map whose matrix representation with respect to the standard basis $(e_{j})$ is $\beta I$ (see the discussion of [GMP13] §3.1). On the other hand, multiplication of a fixed vector $v\in{\cal V}$ by scalars, i.e., $[v]:\mathbb{H}\to{\cal V}:\alpha\mapsto v\alpha$ is an $\mathbb{H}$-linear map:
$[v](\beta_{1}\alpha_{1}+\beta_{2}\alpha_{2})=v(\beta_{1}\alpha_{1}+\beta_{2}\alpha_{2})=(v\beta_{1})\alpha_{1}+(v\beta_{2})\alpha_{2}=([v]\beta_{1})\alpha_{1}+([v]\beta_{2})\alpha_{2}.$
Its adjoint $[v]^{*}:{\cal V}\to\mathbb{H}$ is given by $[v]^{*}=\langle\cdot,v\rangle$, since
$\langle[v]^{*}w,\alpha\rangle=\langle w,[v]\alpha\rangle=\langle w,v\alpha\rangle=\overline{\alpha}\langle w,v\rangle=\langle\alpha,\langle w,v\rangle\rangle.$
@@
$V^{*}:{\cal V}\to\mathbb{H}^{n}:v\mapsto(\langle v,v_{j}\rangle)_{j=1}^{n}.$
@@ 2 Tight frames
A frame for a Hilbert space ${\cal H}$ is a sequence of vectors $(v_{j})$ satisfying the condition
$A\|v\|^{2}\leq\sum_{j}|\langle v,v_{j}\rangle|^{2}\leq B\|v\|^{2},\qquad\forall v\in{\cal H},$ (2.5)
where $A,B>0$ are constants, with the case $A=B$ giving a tight frame. From this, a “frame expansion” follows, which for tight frames takes the particularly simple form
$v={1\over A}\sum_{j}v_{j}\langle v,v_{j}\rangle,\qquad\forall v\in{\cal H}.$
A prominent early example of the use of such “generalised orthonormal bases” is in the theory of wavelets. Recently, frames have been considered for quaternionic Hilbert space, see, e.g., [KTS17] and [VSSS20]. These papers deal primarily with the the frame operator and the construction of dual frames. Here we consider tight frames (where the dual frame is the frame itself) with a particular emphasis on the classification and construction of such frames. This is related to earlier work of Hoggar [Hog77], [Hog82] and others, which implicitly considers tight frames over quaternionic (and even octonionic) Hilbert spaces.
@@ 2.1 Tight frames defined and unitary equivalence
We will say that a sequence of vectors with synthesis map $V=[v_{1},\ldots,v_{n}]$ is a tight frame for a (finite-dimensional) quaternionic Hilbert space ${\cal H}$ if it satisfies (2.5), where $A=B$. It is said to be normalised if $A=1$, which can be achieved by multiplying the vectors by a suitable positive scalar. The frame operator (for a sequence of vectors) is $S=VV^{*}$ and the Gramian (matrix) is $G=V^{*}V$.
A linear map $U$ is unitary if it preserves angles, i.e., $\langle Uv,Uw\rangle=\langle v,w\rangle$, $\forall v,w$, or, equivalently $U^{*}U=I$. Unitary maps can be defined in the same way on quaternionic Hilbert spaces. If $V=[v_{1},\ldots,v_{n}]$ is a frame for a quaternionic Hilbert space, then so is any unitary image $UV=[Uv_{1},\ldots,Uv_{n}]$, and these frames have the same Gramian since $(UV)^{*}UV=V^{*}U^{*}UV=V^{*}V$, and we say that they are unitarily equivalent. Tight frames are studied up to unitary equivalence (which is an equivalence relation) and multiplication by a nonzero scalar.
@@
$\langle v,w\rangle={1\over 4}\sum_{r=0}^{m-1}\Bigl{(}\|v+wi_{r}\|^{2}-\|v-wi_{r}\|^{2}\Bigr{)}i_{r},$
[…]
@@
$\|v\pm wi_{r}\|^{2}$ $=\langle v,v\rangle+\langle\pm wi_{r},\pm wi_{r}\rangle+\langle v,\pm wi_{r}\rangle+\langle\pm wi_{r},v\rangle$
@@
$\|v+wi_{r}\|^{2}-\|v-wi_{r}\|^{2}=2\bigl{(}\overline{i_{r}}\langle v,w\rangle+\overline{\langle v,w\rangle}i_{r}\bigr{)}=4(\langle v,w\rangle)_{r},$
@@
Let $V=[v_{1},\ldots,v_{n}]$ be sequence of vectors in a $d$-dimensional (right) quaternionic Hilbert space ${\cal H}$, such as $\mathbb{H}^{d}$. Then the following are equivalent
@@
    $V$ is a normalised tight frame for ${\cal H}$, i.e.,
    $\|v\|^{2}=\sum_{j}|\langle v,v_{j}\rangle|^{2},\qquad\forall v\in{\cal H}.$
@@
    $v=\sum_{j}v_{j}\langle v,v_{j}\rangle,\qquad\forall v\in{\cal H}.$
@@
    $\langle v,w\rangle=\sum_{j}\langle v_{j},w\rangle\langle v,v_{j}\rangle,\qquad\forall v,w\in{\cal H}.$
@@
    The Gramian $P=V^{*}V$ is a rank $d$ orthogonal projection, i.e., $P^{2}=P$, $P^{*}=P$.
Proof: The implications (ii)$\Longrightarrow$(iii)$\Longrightarrow$(i) follow by taking the inner product with $w$ and then letting $w=v$, respectively. Suppose that (i) holds. By Lemma 2.1, we have
$4(\langle v,w\rangle)_{r}$ $=\|v+wi_{r}\|^{2}-\|v-wi_{r}\|^{2}=\sum_{j}\Bigl{(}|\langle v+wi_{r},v_{j}\rangle|^{2}-|\langle v-wi_{r},v_{j}\rangle|^{2}\Bigr{)}$
$=\sum_{j}\Bigl{(}2\overline{i_{r}}\langle v_{j},w\rangle\langle v,v_{j}\rangle+2\langle v_{j},v\rangle\langle w,v_{j}\rangle i_{r}\Bigr{)}=4\sum_{j}(\langle v_{j},w\rangle\langle v,v_{j}\rangle)_{r}.$
@@
$\langle v,w\rangle=\sum_{j}\langle f_{j},w\rangle\langle v,f_{j}\rangle=\langle\sum_{j}f_{j}\langle v,f_{j}\rangle,w\rangle\hskip 10.00002pt\Longrightarrow\hskip 10.00002ptv=\sum_{j}f_{j}\langle v,f_{j}\rangle,$
@@
We now show (iii)$\iff$(iv). We observe that by construction $P=(\langle v_{k},v_{j}\rangle)_{j,k}$ is Hermitian. The condition $P^{2}=P$ can be written entrywise as
$\langle v_{k},v_{j}\rangle=P_{jk}=\sum_{\ell}P_{j\ell}P_{\ell k}=\sum_{\ell}\langle v_{\ell},v_{j}\rangle\langle v_{k},v_{\ell}\rangle,$
which is the Plancherel identity for $v=v_{k}$ and $w=v_{j}$. The implications then follow by extending the Plancherel identity (using linearity and symmetry of the inner product), and calculating $\mathop{\rm rank}\nolimits(P)=\mathop{\rm trace}\nolimits(P)=\mathop{\rm Re}\nolimits(\mathop{\rm trace}\nolimits(VV^{*}))=d$, by (ii).
For ease of presentation, we will now consider $\mathbb{H}^{d}$, rather than saying let ${\cal H}$ be a quaternionic Hilbert space of dimension $d$. We also write $\mathbb{F}^{d}$, with $\mathbb{F}=\mathbb{R},\mathbb{C},\mathbb{H}$. The following characterisation extends the real and complex cases (see [Wal18] Theorem 2.1).
@@
$\langle v_{k},v_{j}\rangle=\langle Pe_{k},Pe_{k}\rangle=\langle e_{k},Pe_{j}\rangle=\overline{P_{kj}}=P_{jk},$
@@
$|\langle v_{j},v_{k}\rangle|^{2}=\lambda=c^{2}=\cos\theta,\qquad\forall j\neq k.$ (2.7)
@@ 2.2 The variational characterisation of tight frames
We now seek to extend the variational characterisation for tight frames [BF03], [Wal03]. This is most easily proved over $\mathbb{C}$ from the spectral decomposition of the frame operator using the formula $\mathop{\rm trace}\nolimits(AB)=\mathop{\rm trace}\nolimits(BA)$ (see [Wal18], Theorem 6.1). This trace formula no longer holds over the quaternions, even for $1\times 1$ matrices. Instead, we will use the fact
@@
$Av=v\lambda\hskip 10.00002pt\Longrightarrow\hskip 10.00002ptA(v\alpha)=(v\alpha)\alpha^{-1}\lambda\alpha,$
@@
The formulas for $\mathop{\rm trace}\nolimits(G)$ and $\mathop{\rm trace}\nolimits(G^{2})$ given on the left hand side of (2.9) are easily calculated from $(G)_{jk}=\langle v_{k},v_{j}\rangle$.
@@
Let $v_{1},\ldots,v_{n}$ be vectors in $\mathbb{F}^{d}$, which are not all zero. Then
@@
$S=U(AI)U^{*}=AI\hskip 10.00002pt\Longleftrightarrow\hskip 10.00002pt\hbox{$(v_{j})$ is a tight frame for $\mathbb{F}^{d}$.}$

### v3

@@ 1 Introduction
[…]
should hold. However, calculations of [CKM16], suggests that this fails for $d=4$, and the analogous situation for the octonians is much worse. Thus, it seems that equiangular lines in $\mathbb{C}^{d}$ may be a high point for satisfying the estimates on the maximal number of equiangular lines, with real and quaternionic equiangular lines rarely meeting the bound (“filling up all the space”) due to algebraic limitations of the field involved, i.e., $\mathbb{R}$ not being algebraically closed and $\mathbb{H}$ not being commutative. Still, there is much interest in the maximal sets of equiangular lines in $\mathbb{H}^{d}$, and for those in $\mathbb{R}^{d}$ (which have been studied for over half a century). Therefore, I present the following conjecture, which can play the role of Zauner’s conjecture for the theory of quaternionic equiangular lines:
@@ 1.1 Inner products over the quaternions
Since the multiplication is not commutative, we must distinguish between left and right vector spaces (modules) over $\mathbb{H}$. Since we wish to appropriate much of matrix theory, we take our vector spaces to be right $\mathbb{H}$-vector spaces. Thus $\mathbb{H}$-linear maps $L$ have the form
@@
where order of multiplication in $a_{j\ell}b_{\ell k}$ cannot be reversed (see [Zha97]). For those who may have noticed, I apologise for using $j$ and $k$ above as indices for matrix entries, and elsewhere as quaternion units (as is often done with the complex unit $i$).
@@
$\overline{ab}=\overline{b}\,\overline{a},\qquad a,b\in\mathbb{H}\quad\Longrightarrow\quad(AB)^{*}=B^{*}A^{*}\quad\hbox{(for matrices over $\mathbb{H}$)}.$
@@
$\langle v,w\rangle:=v^{*}w=\sum_{j}\overline{v_{j}}w_{j},\qquad v,w\in\mathbb{H}^{d}.$ (1.2)
@@
In light of the noncommutativity of the quaternions, we note that scalars come outside an inner product (as we have defined it) as follows
@@
$\langle v,w\rangle=0\quad\Longleftrightarrow\quad\langle w,v\rangle=\overline{\langle v,w\rangle}=0.$
@@
If $T$ and $T^{*}$ are represented as matrices $[T]$ and $T^{*}]$ with respect to orthonormal bases $(v_{j})$ and $(w_{k})$, so that $v=\sum_{j}v_{j}\langle v_{j},v\rangle$, $\forall v\in{\cal V}$, and $w=\sum_{k}w_{k}\langle w_{k},w\rangle$, $\forall w\in{\cal W}$, then
$T]_{jk}=\langle w_{j},Tv_{k}\rangle=\langle T^{*}w_{j},v_{k}\rangle=\overline{\langle v_{k},T^{*}w_{j}\rangle}=\overline{[T^{*}]_{kj}},$
and hence the matrix $T^{*}]$ is the conjugate transpose of the matrix $[T]$. For this reason, it is often assumed that the inner product is the standard inner product on $\mathbb{H}^{d}$, and all calculations are done with matrices, with $A^{*}$ defined to be the conjugate transpose (or Hermitian transpose) of the matrix $A$, as is the case in [Rod14]. The adjoint and Hermitian transpose satisfy some (but not all) of the usual properties, including
@@
$(A^{*})^{-1}=(A^{-1})^{*}\quad\hbox{(for $A$ invertible)}.$
Note that the transpose does not satisfy $(AB)^{T}=B^{T}A^{T}$ (since $\mathbb{H}$ is not commutative). It can be shown that if $AB=I$ for square matrices over $\mathbb{H}$ then $BA=I$, and so a right inverse exists for $A$ if and only if a left inverse exists, and these inverses are equal (and denoted by $A^{-1}$).
@@
$R_{\beta}(v\alpha)=(v\alpha)\beta=v(\alpha\beta)\neq v(\beta\alpha)=(v\beta)\alpha=(R_{\beta}v)\alpha\qquad\hbox{(in general)}.$
Left multiplication of $\mathbb{H}^{d}$ by $\beta$ defines an $\mathbb{H}$-linear map $L_{\beta}:\mathbb{H}^{d}\to\mathbb{H}^{d}:v\mapsto\beta v$, but this is dependent on a choice of basis: it is the linear map which maps $e_{j}\mapsto e_{j}\beta$, i.e., the linear map whose matrix representation with respect to the standard basis $(e_{j})$ is $\beta I$ (see the discussion of [GMP13] §3.1). On the other hand, multiplication of a fixed vector $v\in{\cal V}$ by scalars, i.e., $v]:\mathbb{H}\to{\cal V}:\alpha\mapsto v\alpha$ is an $\mathbb{H}$-linear map:
$v](\beta_{1}\alpha_{1}+\beta_{2}\alpha_{2})=v(\beta_{1}\alpha_{1}+\beta_{2}\alpha_{2})=(v\beta_{1})\alpha_{1}+(v\beta_{2})\alpha_{2}=([v]\beta_{1})\alpha_{1}+([v]\beta_{2})\alpha_{2}.$
Its adjoint $v]^{*}:{\cal V}\to\mathbb{H}$ is given by $v]^{*}=\langle v,\cdot\rangle$, since
$\langle\alpha,[v]^{*}w\rangle=\langle[v]\alpha,w\rangle=\langle v\alpha,w\rangle=\overline{\alpha}\langle v,w\rangle=\overline{\alpha}\langle 1,\langle v,w\rangle\rangle=\langle\alpha,\langle v,w\rangle\rangle.$
@@
$V^{*}:{\cal V}\to\mathbb{H}^{n}:v\mapsto(\langle v_{j},v\rangle)_{j=1}^{n}.$
@@ 2 Tight frames
[…]
$A\|v\|^{2}\leq\sum_{j}|\langle v_{j},v\rangle|^{2}\leq B\|v\|^{2},\qquad\forall v\in{\cal H},$ (2.5)
where $A,B>0$ are constants. From this, a “frame expansion” follows, which takes a particularly simple form when $A=B$, i.e.,
$v={1\over A}\sum_{j}v_{j}\langle v_{j},v\rangle,\qquad\forall v\in{\cal H}.$
A prominent early example of the use of such “generalised orthonormal bases” is in the theory of wavelets. Recently, frames have been considered for quaternionic Hilbert space, see, e.g., [KTS17], [VSSS20] (which deal primarily with the frame operator and the construction of dual frames), [KPS21] (equiangular lines and Hadamard matrices), and [NS22] which considers the connectedness of the algebraic variety of quaternionic frames with given norms. Here we consider tight frames (where the dual frame is the frame itself) with a particular emphasis on the classification and construction of such frames. This is related to earlier work of Hoggar [Hog77], [Hog82] and others, which implicitly considers tight frames over quaternionic (and even octonionic) Hilbert spaces.
@@ 2.1 Tight frames defined and unitary equivalence
[…]
@@
$\langle v,w\rangle={1\over 4}\sum_{r=0}^{m-1}\Bigl{(}\|vi_{r}+w\|^{2}-\|vi_{r}-w\|^{2}\Bigr{)}i_{r},$
where $m=\dim_{\mathbb{R}}(\mathbb{F})$, $(i_{0},i_{1},i_{2},i_{3})=(1,i,j,k)$, and $\langle\cdot,\cdot\rangle$ is linear in the second variable.
@@
$\|vi_{r}\pm w\|^{2}$ $=\langle vi_{r},vi_{r}\rangle+\langle\pm w,\pm w\rangle+\langle vi_{r},\pm w\rangle+\langle\pm w,vi_{r}\rangle$
@@
$\|vi_{r}+w\|^{2}-\|vi_{r}-w\|^{2}=2\bigl{(}\overline{i_{r}}\langle v,w\rangle+\overline{\langle v,w\rangle}i_{r}\bigr{)}=4(\langle v,w\rangle)_{r},$
@@
[…]
@@
[…]
    $\|v\|^{2}=\sum_{j}|\langle v_{j},v\rangle|^{2},\qquad\forall v\in{\cal H}.$
@@
    $v=\sum_{j}v_{j}\langle v_{j},v\rangle,\qquad\forall v\in{\cal H}.$
@@
    $\langle v,w\rangle=\sum_{j}\langle v,v_{j}\rangle\langle v_{j},w\rangle,\qquad\forall v,w\in{\cal H}.$
@@
    The Gramian $P=V^{*}V$ is a rank $d$ orthogonal projection, i.e., $^{2}=P, P=P$$^{*}=P$.
[…]
$4(\langle v,w\rangle)_{r}$ $=\|vi_{r}+w\|^{2}-\|vi_{r}-w\|^{2}=\sum_{j}\Bigl{(}|\langle v_{j},vi_{r}+w\rangle|^{2}-|\langle v_{j},vi_{r}-w\rangle|^{2}\Bigr{)}$
$=\sum_{j}\Bigl{(}2\overline{i_{r}}\langle v,v_{j}\rangle\langle v_{j},w\rangle+2\langle w,v_{j}\rangle\langle v_{j},v\rangle i_{r}\Bigr{)}=4\sum_{j}(\langle v,v_{j}\rangle\langle v_{j},w\rangle)_{r}.$
@@
$\langle v,w\rangle=\sum_{j}\overline{\langle v_{j},v\rangle}\langle v_{j},w\rangle=\langle\sum_{j}v_{j}\langle v_{j},v\rangle,w\rangle\quad\Longrightarrow\quad v=\sum_{j}v_{j}\langle v_{j},v\rangle,$
@@
[…]
$\langle v_{j},v_{k}\rangle=P_{jk}=\sum_{\ell}P_{j\ell}P_{\ell k}=\sum_{\ell}\langle v_{j},v_{\ell}\rangle\langle v_{\ell},v_{k}\rangle,$
which is the Plancherel identity for $v=v_{j}$ and $w=v_{k}$. The implications then follow by extending the Plancherel identity (using linearity and symmetry of the inner product), and calculating $\mathop{\rm rank}\nolimits(P)=\mathop{\rm trace}\nolimits(P)=\mathop{\rm Re}\nolimits(\mathop{\rm trace}\nolimits(VV^{*}))=d$, by (ii).
[…]
@@
$\langle v_{j},v_{k}\rangle=\langle Pe_{j},Pe_{k}\rangle=\langle e_{j},Pe_{k}\rangle=P_{jk},$
@@
$|\langle v_{j},v_{k}\rangle|^{2}=\lambda=c^{2}=(\cos\theta)^{2},\qquad\forall j\neq k.$ (2.7)
@@ 2.2 The variational characterisation of tight frames
We now seek to extend the variational characterisation for tight frames [BF03], [Wal03]. For $\mathbb{C}^{d}$, this is most easily and transparently proved from the spectral decomposition of the frame operator using $\mathop{\rm trace}\nolimits(AB)=\mathop{\rm trace}\nolimits(BA)$ (see [Wal18], Theorem 6.1). This trace formula no longer holds over the quaternions, even for $1\times 1$ matrices. Instead, we will use the fact
@@
$Av=v\lambda\quad\Longrightarrow\quad A(v\alpha)=(v\alpha)\alpha^{-1}\lambda\alpha,$
@@
The formulas for $\mathop{\rm trace}\nolimits(G)$ and $\mathop{\rm trace}\nolimits(G^{2})$ given on the left hand side of (2.9) are easily calculated from $(G)_{jk}=\langle v_{j},v_{k}\rangle$.
@@
(Variational characterisation) Let $v_{1},\ldots,v_{n}$ be vectors in $\mathbb{F}^{d}$, which are not all zero. Then
@@
$S=U(AI)U^{*}=AI\quad\Longleftrightarrow\quad\hbox{$(v_{j})$ is a tight frame for $\mathbb{F}^{d}$.}$